I don't know if this is the right category, but here is my problem:
The equation of the line containing one side of a parallelogram is 3x + 2y = 8.
The opposite side contains the point (0,-7).
Which of these is the equation of the line that contains the opposite side?
A y = 2/3x-7
B y = -3/2x+7
C y = 2/3x+7
D y = -3/2x-7
Please help out.

Answer by stanbon(75887)

The equation of the line containing one side of a parallelogram is 3x + 2y = 8.
The opposite side contains the point (0,-7).
Which of these is the equation of the line that contains the opposite side?
A y = 2/3x-7
B y = -3/2x+7
C y = 2/3x+7
D y = -3/2x-7
-------------------------
Find the slope of the given line:
3x+2y=8
y = (-3/2)x + 4
slope is -3/2
---------------------
Since the figure is a parallelogram, the opposite side is parallel and
has the same slope. So you want the equation of the line that has
slope = -3/2 and passes thru the point (0,-7)
--------------
Form of the equation is y=mx+b and you know y=-7 when x=0 and you know m=-3/2
So, solve for "b":
-7 = (-3/2)*0 + b
b = -7
-------------------
EQUATION:
y = (-3/2)x - 7
=======================
Cheers,
Stan H.